OpenStudy (ccchristian1):

How many solutions does the equation 4x + 8 = − 2 − 2 + 5x have? Two One Zero Infinitely many

OpenStudy (ccchristian1):

i think this one is d

pooja195 (pooja195):

No

OpenStudy (ccchristian1):

zero?

OpenStudy (anonymous):

no

pooja195 (pooja195):

We need to simplify 4x+8=−2−2+5x 4x+8=−2+−2+5x 4x+8=(5x)+(−2+−2)(Combine Like Terms) 4x+8=5x+−4 4x+8=5x−4 Step 2: Subtract 5x from both sides. 4x+8−5x=5x−4−5x −x+8=−4 Step 3: Subtract 8 from both sides. −x+8−8=−4−8 −x=−12

pooja195 (pooja195):

Finish it off :)

OpenStudy (anonymous):

actually it is positive x

OpenStudy (ccchristian1):

so one solution ?

pooja195 (pooja195):

Yes

pooja195 (pooja195):

@unkownbeast911 i subtracted X

OpenStudy (anonymous):

actually all of it is positive

OpenStudy (ccchristian1):

so is it 1 solution?

pooja195 (pooja195):

Either way it still has one solution .-.

OpenStudy (ccchristian1):

okay:)

OpenStudy (anonymous):

true